# Dual-Source SPIR over a noiseless MAC without Data Replication or Shared Randomness

### v1

@@ IV-C3 Servers’ mutual privacy
$=I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}})+I(F_{1,\mathcal{L}_{1}};F_{2,\mathcal{L}_{2}}|\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K})$
=$\stackrel{{\scriptstyle(a)}}{{=}}I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}})$
@@
$=I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};\bar{C}_{2,1:K})+I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})$
=$\stackrel{{\scriptstyle(b)}}{{=}}I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};\bar{C}_{2,1:K})$
$\xrightarrow{n\to\infty}0,$
@@
    holds because $I(F_{1,\mathcal{L}_{1}};F_{2,\mathcal{L}_{2}}|\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K})\leqslant I(F_{1,\mathcal{L}_{1}}\bar{C}_{1,1:K}X^{n}_{1,1:K};F_{2,\mathcal{L}_{2}}|\mathbf{A}_{1:K})=0$;
@@
    holds because $I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})=I(\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})+I(\bar{C}_{1,1:K}X^{n}_{1,1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K}\mathbf{A}_{1:K})\leqslant I(\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})+I(\bar{C}_{1,1:K}X^{n}_{1,1:K};F_{2,\mathcal{L}_{2}}\bar{C}_{2,1:K}|\mathbf{A}_{1:K})=I(\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})=0$;
  3. and the limit holds by (12).
@@
$I(F_{2,\mathcal{L}_{2}}X^{n}_{2,1:K}\mathbf{A}_{1:K};F_{1,\mathcal{L}_{1}})\xrightarrow{n\to\infty}0.$
@@ V Converse part of Theorem 1
$\mathcal{C}_{\textup{SPIR}^{2}}(L_{1},L_{2})\subseteq\left\{(R_{1},R_{2}):(L_{1}-1)R_{1}+(L_{2}-1)R_{2}\leqslant\max_{p_{X_{1}}p_{X_{2}}}H(X_{1}X_{2}|Y)\right\}.$
@@
It is sufficient to prove (13), as the proof of (14) can be obtained by exchanging the roles of the servers. Define for $t\in\llbracket 1,n\rrbracket$, $i\in\llbracket 1,r_{t}\rrbracket$, $A_{1,t,1:i}\triangleq(M_{0,1,t}(j),M_{1,t}(j))_{j\in\llbracket 1,i\rrbracket}$ the messages sent and received by Server 1 between the first and the $i$-th communication exchanges with the client that happen after the $t$-th channel use. Let $A_{1}^{t}\triangleq(A_{1,j,1:r_{j}})_{j\in\llbracket 1,t\rrbracket}$ be all the messages sent and received by Server 1 before the $t+1$-th channel use. Let $t\in\llbracket 1,n\rrbracket$ and $j\in\llbracket 1,r_{t}\rrbracket$. We have
$I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t}U_{0}|X_{1}^{t}A_{1}^{t-1}A_{1,t,1:j}Z_{1}Z_{2})$ (15)
$=I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t}U_{0}|X_{1}^{t}A_{1}^{t-1}A_{1,t,1:(j-1)}M_{0,1,t}(j)M_{1,t}(j)Z_{1}Z_{2})$
$\leqslant I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t}U_{0}M_{0,1,t}(j)|X_{1}^{t}A_{1}^{t-1}A_{1,t,1:(j-1)}M_{1,t}(j)Z_{1}Z_{2})$
=$\stackrel{{\scriptstyle(a)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t}U_{0}|X_{1}^{t}A_{1}^{t-1}A_{1,t,1:(j-1)}M_{1,t}(j)Z_{1}Z_{2})$
$\leqslant I(F_{1,\mathcal{L}_{1}}U_{1}M_{1,t}(j);Y^{t}U_{0}|X_{1}^{t}A_{1}^{t-1}A_{1,t,1:(j-1)}Z_{1}Z_{2})$
=$\stackrel{{\scriptstyle(b)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t}U_{0}|X_{1}^{t}A_{1}^{t-1}A_{1,t,1:(j-1)}Z_{1}Z_{2})$ (16)
$\stackrel{{\scriptstyle(c)}}{{\leqslant}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t}U_{0}|X_{1}^{t}A_{1}^{t-1}Z_{1}Z_{2})$
=$\stackrel{{\scriptstyle(d)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t-1}U_{0}|X_{1}^{t}A_{1}^{t-1}Z_{1}Z_{2})$
$\leqslant I(F_{1,\mathcal{L}_{1}}U_{1}(X_{1})_{t};Y^{t-1}U_{0}|X_{1}^{t-1}A_{1}^{t-1}Z_{1}Z_{2})$
=$\stackrel{{\scriptstyle(e)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t-1}U_{0}|X_{1}^{t-1}A_{1}^{t-1}Z_{1}Z_{2}),$ (17)
@@
    holds by the chain rule and because by definition $M_{0,1,t}(j)$ is a function of $(Z_{1},Z_{2},U_{0},Y^{t},A_{1}^{t-1},A_{1,t,1:(j-1)},M_{1,t}(j))$;
@@
    holds by the chain rule and because $M_{1,t}(j)$ is a function of $(F_{1,\mathcal{L}_{1}},U_{1},A_{1}^{t-1},A_{1,t,1:(j-1)})$;
@@
    holds by repeating $j-1$ times the steps between Equations (15) and (16);
@@
    holds because $(F_{1,\mathcal{L}_{1}},U_{1})-(Y^{t-1},U_{0},X_{1}^{t},A_{1}^{t-1},Z_{1},Z_{2})-Y_{t}$ forms a Markov chain;
@@
    holds by the chain rule and because $(X_{1})_{t}$ is a function of $(F_{1,\mathcal{L}_{1}},U_{1},A_{1}^{t-1})$.
Then, for any $t\in\llbracket 1,n\rrbracket$, we have
$I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t}U_{0}|X_{1}^{t}A_{1}^{t}Z_{1}Z_{2})$ $\stackrel{{\scriptstyle(a)}}{{\leqslant}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{t-1}U_{0}|X_{1}^{t-1}A_{1}^{t-1}Z_{1}Z_{2})$ (18)
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}I(F_{1,\mathcal{L}_{1}}U_{1};U_{0}|Z_{1}Z_{2})$
$=0,$ (19)
where
  1. (a)
    holds by taking $j=r_{t}$ in (17);
  2. (b)
    holds by using $t-1$ times (18).
@@
where the last equality holds by taking $t=n$ in (19). ∎
@@
Lemma 5 ([9, Lemma 3]).
@@
$|H(X|Y,Z=z_{1})-H(X|Y,Z=z_{2})|\leqslant 1+3\log|\mathcal{X}|\sqrt{\frac{(p+q)\ln 2}{2pq}I(XY;Z)}.$
@@
$\stackrel{{\scriptstyle(a)}}{{\leqslant}}H(F_{1,\mathcal{L}_{1}\backslash\{z_{1}\}}|X_{1}^{n}\mathbf{A}_{1},Z_{1}=\gamma_{1},Z_{2}={z}_{2})+6nR_{1}\sqrt{\ln 2}\sqrt{I(F_{1,\mathcal{L}_{1}\backslash\{z_{1}\}}X_{1}^{n}\mathbf{A}_{1};Z_{1}Z_{2})}+1$
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}H(F_{1,\mathcal{L}_{1}\backslash\{z_{1}\}}|X_{1}^{n}\mathbf{A}_{1},Z_{1}=\gamma_{1},Z_{2}={z}_{2})+o(n)$
$\stackrel{{\scriptstyle(c)}}{{\leqslant}}H(F_{1,\mathcal{L}_{1}\backslash\{z_{1},\gamma_{1}\}}|X_{1}^{n}\mathbf{A}_{1},Z_{1}=\gamma_{1},Z_{2}={z}_{2})+H(F_{1,\gamma_{1}}|X_{1}^{n}\mathbf{A}_{1},Z_{1}=\gamma_{1},Z_{2}={z}_{2})+o(n)$
@@
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{i=1}^{L_{1}-1}H(F_{1,\gamma_{i}}|Y^{n}U_{0}X_{1}^{n}\mathbf{A}_{1},Z_{1}=\gamma_{i},Z_{2}={z}_{2})+o(n)$
@@
$=\sum_{z_{1},z_{2}}\mathbb{P}[(Z_{1},Z_{2})=(z_{1},z_{2})]H(F_{1,\mathcal{L}_{1}\backslash\{Z_{1}\}}|X_{1}^{n}\mathbf{A}_{1},Z_{1}=z_{1},Z_{2}=z_{2})$
@@
$=H(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}}|Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})+I(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}};Y^{n}\mathbf{A}|{Z}_{1}{Z}_{2})$
$\leqslant H(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}}|Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})+I(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}};Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})$
$\stackrel{{\scriptstyle(a)}}{{\leqslant}}H(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}}|Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})+o(n)$
@@
$=H(X_{1}^{n}X_{2}^{n}|Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})+H(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}}|X_{1}^{n}X_{2}^{n}Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})+o(n)$
$\stackrel{{\scriptstyle(b)}}{{\leqslant}}H(X_{1}^{n}X_{2}^{n}|Y^{n})+H(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}|X_{1}^{n}\mathbf{A}{Z}_{1}{Z}_{2})+H(F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}}|X_{2}^{n}\mathbf{A}{Z}_{1}{Z}_{2})+o(n)$
@@
$(L_{1}-1)nR_{1}+(L_{2}-1)nR_{2}$ =$\stackrel{{\scriptstyle(a)}}{{=}}H(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}}|{Z}_{1}{Z}_{2})$
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}H(X_{1}|Y)$
=$\stackrel{{\scriptstyle(b)}}{{=}}H(X_{1}|Y=1)\mathbb{P}[Y=1],$
@@
$f:]0,1[\times]0,1[\to\mathbb{R}_{+},(p_{1},p_{2})\mapsto-\bar{p}_{1}p_{2}\log\frac{\bar{p}_{1}p_{2}}{p_{1}\bar{p}_{2}+\bar{p}_{1}p_{2}}-p_{1}\bar{p}_{2}\log\frac{p_{1}\bar{p}_{2}}{p_{1}\bar{p}_{2}+\bar{p}_{1}p_{2}},$
@@
Step 2. We now prove that a necessary and sufficient condition for the maximum in (24) to be achieved is $p_{1}+p_{2}=1$. After computations, we have
$\frac{\partial f}{\partial p_{1}}(p_{1},p_{2})+\frac{\partial f}{\partial p_{2}}(p_{1},p_{2})=(p_{1}+p_{2}-1)\log\frac{p_{1}\bar{p}_{1}p_{2}\bar{p}_{2}}{(p_{1}\bar{p}_{2}+\bar{p}_{1}p_{2})^{2}}.$ (25)
@@
Step 4. We now show that $\max_{p_{1}\in[0,1/2]}g(p_{1})=g(1/2)$. Specifically, it is sufficient to show that $g$ is non-decreasing over ]0,1/2]]0,1/2]] 0 , 1 / 2 ] by showing that the derivative $g^{\prime}$ of $g$ is non-negative over ]0,1/2]]0,1/2]] 0 , 1 / 2 ]. We have for $x\in]0,1/2]$, $g^{\prime}(x)=2(1-x)\log\frac{(1-x)^{2}}{(1-x)^{2}+x^{2}}-2x\log\frac{x^{2}}{(1-x)^{2}+x^{2}}$. Next, define $k:[1,+\infty[\to\mathbb{R},y\mapsto\log(1+y^{2})-y\log(1+y^{-2})$ such that, by the change of variable $y\leftarrow(x^{-1}-1)$, we have $\forall x\in]0,1/2],g^{\prime}(x)\geqslant 0\Leftrightarrow\forall y\in[1,+\infty[,k(y)\geqslant 0$. Since $k(1)=0$, to show that $\forall y\in[1,+\infty[,k(y)\geqslant 0$, it is sufficient to show that $k$ is non-decreasing over [1,+[[1,$\infty$[[ 1 , + ∞ [ by showing that the derivative $k^{\prime}$ of $k$ is non-negative over [1,+[[1,$\infty$[[ 1 , + ∞ [. We have for $y\in[1,+\infty[$, $k^{\prime}(y)=\frac{2}{\ln 2}\frac{1+y}{1+y^{2}}-\log(1+y^{-2})$. Since $k^{\prime}(1)=\frac{2-\ln 2}{\ln 2}>0$ and $\lim_{y\to+\infty}k^{\prime}(y)=0$, to show that $k^{\prime}$ is non-negative over [1,+[[1,$\infty$[[ 1 , + ∞ [, it is sufficient to show that $k^{\prime}$ is non-increasing over [1,+[[1,$\infty$[[ 1 , + ∞ [, which is true since the derivative $k^{{}^{\prime\prime}}$ of $k^{\prime}$ satisfies for $y\in[1,+\infty[$, $k^{{}^{\prime\prime}}(y)=\frac{2(1-y)(1+y)^{2}}{y(1+y^{2})^{2}\ln 2}\leqslant 0$.

### v2

@@ IV-C3 Servers’ mutual privacy
$=I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}})$
$\phantom{-}+I(F_{1,\mathcal{L}_{1}};F_{2,\mathcal{L}_{2}}|\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K})$
$\stackrel{{\scriptstyle(a)}}{{=}}I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}})$
@@
$=I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};\bar{C}_{2,1:K})$
$\phantom{-}+I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})$
$\stackrel{{\scriptstyle(b)}}{{=}}I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};\bar{C}_{2,1:K})$
$\stackrel{{\scriptstyle(c)}}{{=}}0,$
@@
    holds because
    $I(F_{1,\mathcal{L}_{1}};F_{2,\mathcal{L}_{2}}|\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K})$
    $\leqslant I(F_{1,\mathcal{L}_{1}}\bar{C}_{1,1:K}X^{n}_{1,1:K};F_{2,\mathcal{L}_{2}}|\mathbf{A}_{1:K})$
    $=0;$
@@
    holds because
    $I(\bar{C}_{1,1:K}X^{n}_{1,1:K}\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})$
    $=I(\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})$
    $\phantom{-}I(\bar{C}_{1,1:K}X^{n}_{1,1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K}\mathbf{A}_{1:K})$
    $\leqslant I(\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})$
    $\phantom{-}+I(\bar{C}_{1,1:K}X^{n}_{1,1:K};F_{2,\mathcal{L}_{2}}\bar{C}_{2,1:K}|\mathbf{A}_{1:K})$
    $=I(\mathbf{A}_{1:K};F_{2,\mathcal{L}_{2}}|\bar{C}_{2,1:K})=0;$
  3. (c)
    holds by (14).
@@
$I(F_{2,\mathcal{L}_{2}}X^{n}_{2,1:K}\mathbf{A}_{1:K};F_{1,\mathcal{L}_{1}})=0.$
@@ IV-C4 Rates
Remark.
The download cost from Server $i\in\{1,2\}$ is $2nR_{i}(L_{i}-1)$ bits, i.e., $2(L_{i}-1)$ bits per bit of downloaded file. For Server $i\in\{1,2\}$, the total number of bitwise XOR operations is $2nR_{i}(2L_{i}-3)$, i.e., $2(2L_{i}-3)$ bitwise XOR operations per bit of downloaded file.
@@ V Converse part of Theorem 1
$\mathcal{C}_{\textup{SPIR}^{2}}(L_{1},L_{2})$ $\subseteq\left\{(R_{1},R_{2}):(L_{1}-1)R_{1}+(L_{2}-1)R_{2}\right$.
$\phantom{------l}\left.\leqslant\textstyle\max_{p_{X_{1}}p_{X_{2}}}H(X_{1}X_{2}|Y)\right\}.$
@@
It is sufficient to prove (15), as the proof of (16) can be obtained by exchanging the roles of the servers. We have (19)
$I(F_{1,\mathcal{L}_{1}}U_{1};Y^{n}U_{0}|X_{1}^{n}\mathbf{A}_{1}Z_{1}Z_{2})$ $=I(F_{1,\mathcal{L}_{1}}U_{1};Y^{n}U_{0}|X_{1}^{n}(M_{0,1}(j),M_{1}(j))_{j\in\llbracket 1,r\rrbracket}Z_{1}Z_{2})$ (17)
$\leqslant I(F_{1,\mathcal{L}_{1}}U_{1};Y^{n}U_{0}M_{0,1}(r)|X_{1}^{n}(M_{0,1}(j),M_{1}(j))_{j\in\llbracket 1,r-1\rrbracket}M_{1}(r)Z_{1}Z_{2})$
$\stackrel{{\scriptstyle(a)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{n}U_{0}|X_{1}^{n}(M_{0,1}(j),M_{1}(j))_{j\in\llbracket 1,r-1\rrbracket}M_{1}(r)Z_{1}Z_{2})$
$\leqslant I(F_{1,\mathcal{L}_{1}}U_{1}M_{1}(r);Y^{n}U_{0}|X_{1}^{n}(M_{0,1}(j),M_{1}(j))_{j\in\llbracket 1,r-1\rrbracket}Z_{1}Z_{2})$
$\stackrel{{\scriptstyle(b)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{n}U_{0}|X_{1}^{n}(M_{0,1}(j),M_{1}(j))_{j\in\llbracket 1,r-1\rrbracket}Z_{1}Z_{2})$ (18)
$\stackrel{{\scriptstyle(c)}}{{\leqslant}}I(F_{1,\mathcal{L}_{1}}U_{1};Y^{n}U_{0}|X_{1}^{n}Z_{1}Z_{2})$
$\stackrel{{\scriptstyle(d)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};U_{0}|X_{1}^{n}Z_{1}Z_{2})$
$\leqslant I(F_{1,\mathcal{L}_{1}}U_{1}X_{1}^{n};U_{0}|Z_{1}Z_{2})$
$\stackrel{{\scriptstyle(e)}}{{=}}I(F_{1,\mathcal{L}_{1}}U_{1};U_{0}|Z_{1}Z_{2})$
$=0$ (19)
@@
    holds by the chain rule and because by definition $M_{0,1}(r)$ is a function of $(Z_{1},U_{0},Y^{n},(M_{1}(j))_{j\in\llbracket 1,r\rrbracket})$;
@@
    holds by the chain rule and because $M_{1}(r)$ is a function of $(F_{1,\mathcal{L}_{1}},U_{1},(M_{0,1}(j))_{j\in\llbracket 1,r-1\rrbracket})$;
@@
    holds by repeating r1r-1 times the steps between Equations (17) and (18);
@@
    holds because $(F_{1,\mathcal{L}_{1}},U_{1})-(U_{0},X_{1}^{n},Z_{1},Z_{2})-Y^{n}$ forms a Markov chain;
@@
    holds by the chain rule and because $X_{1}^{n}$ is a function of $(F_{1,\mathcal{L}_{1}},U_{1})$.
@@
where the last equality holds by (19). ∎
@@
Lemma 5 ([ahlswede2013oblivious, Lemma 3]).
@@
$|H(X|Y,Z=z_{1})-H(X|Y,Z=z_{2})|\\ \leqslant 1+3\log|\mathcal{X}|\sqrt{\frac{(p+q)\ln 2}{2pq}I(XY;Z)}.$
@@
[…]
$\phantom{-}+H(F_{1,\gamma_{1}}|X_{1}^{n}\mathbf{A}_{1},Z_{1}=\gamma_{1},Z_{2}={z}_{2})+o(n)$
@@
$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{i=1}^{L_{1}-1}H(F_{1,\gamma_{i}}|Y^{n}U_{0}X_{1}^{n}\mathbf{A}_{1},Z_{1}=\gamma_{i},Z_{2}={z}_{2})+o(n)$
@@
$=\textstyle\sum_{z_{1},z_{2}}\mathbb{P}[(Z_{1},Z_{2})=(z_{1},z_{2})]$
$\phantom{-----}\times H(F_{1,\mathcal{L}_{1}\backslash\{Z_{1}\}}|X_{1}^{n}\mathbf{A}_{1},Z_{1}=z_{1},Z_{2}=z_{2})$
@@
[…]
$\leqslant H(F_{1,\mathcal{L}_{1}\backslash\{{Z}_{1}\}}F_{2,\mathcal{L}_{2}\backslash\{{Z}_{2}\}}|Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})$
[…]
@@
$=H(X_{1}^{n}X_{2}^{n}|Y^{n}\mathbf{A}{Z}_{1}{Z}_{2})$
[…]
@@
$(L_{1}-1)nR_{1}+(L_{2}-1)nR_{2}$
[…]
@@
$\stackrel{{\scriptstyle(a)}}{{=}}H(X_{1}|Y)$
$\stackrel{{\scriptstyle(b)}}{{=}}H(X_{1}|Y=1)\mathbb{P}[Y=1],$
@@
$f:]0,1[\times]0,1[\to\mathbb{R}_{+},$
$(p_{1},p_{2})\mapsto-\bar{p}_{1}p_{2}\log\frac{\bar{p}_{1}p_{2}}{p_{1}\bar{p}_{2}+\bar{p}_{1}p_{2}}-p_{1}\bar{p}_{2}\log\frac{p_{1}\bar{p}_{2}}{p_{1}\bar{p}_{2}+\bar{p}_{1}p_{2}},$
@@
Step 2. We now prove that a necessary condition for the maximum in (24) to be achieved is $p_{1}+p_{2}=1$. After computations, we have
$\frac{\partial f}{\partial p_{1}}(p_{1},p_{2})+\frac{\partial f}{\partial p_{2}}(p_{1},p_{2})\\ =(p_{1}+p_{2}-1)\log\frac{p_{1}\bar{p}_{1}p_{2}\bar{p}_{2}}{(p_{1}\bar{p}_{2}+\bar{p}_{1}p_{2})^{2}}.$ (25)
@@
Step 4. We now show that $\max_{p_{1}\in[0,1/2]}g(p_{1})=g(1/2)$. Specifically, it is sufficient to show that $g$ is non-decreasing over ]0,1/2]]0,1/2] by showing that the derivative $g^{\prime}$ of $g$ is non-negative over ]0,1/2]]0,1/2]. We have for $x\in]0,1/2]$, $g^{\prime}(x)=2(1-x)\log\frac{(1-x)^{2}}{(1-x)^{2}+x^{2}}-2x\log\frac{x^{2}}{(1-x)^{2}+x^{2}}$. Next, define $k:[1,+\infty[\to\mathbb{R},y\mapsto\log(1+y^{2})-y\log(1+y^{-2})$ such that, by the change of variable $y\leftarrow(x^{-1}-1)$, we have $\forall x\in]0,1/2],g^{\prime}(x)\geqslant 0\Leftrightarrow\forall y\in[1,+\infty[,k(y)\geqslant 0$. Since $k(1)=0$, to show that $\forall y\in[1,+\infty[,k(y)\geqslant 0$, it is sufficient to show that $k$ is non-decreasing over $1,+\infty[$ by showing that its derivative is non-negative over $1,+\infty[$. We have for $y\in[1,+\infty[$, $k^{\prime}(y)=\frac{2}{\ln 2}\frac{1+y}{1+y^{2}}-\log(1+y^{-2})$. Since $k^{\prime}(1)=\frac{2-\ln 2}{\ln 2}>0$ and $\lim_{y\to+\infty}k^{\prime}(y)=0$, to show that $k^{\prime}$ is non-negative over $1,+\infty[$, it is sufficient to show that $k^{\prime}$ is non-increasing over $1,+\infty[$, which is true since the derivative $k^{{}^{\prime\prime}}$ of $k^{\prime}$ satisfies for $y\in[1,+\infty[$, $k^{{}^{\prime\prime}}(y)=\frac{2(1-y)(1+y)^{2}}{y(1+y^{2})^{2}\ln 2}\leqslant 0$.